OpenStudy (anonymous):

Help Math!! I will give Medal!! The rectangle below has a total perimeter of 125 cm.

OpenStudy (anonymous):

Write an equation and solve it to find the length of the longer side of the rectangle. 2x + 10 = 125; x = 57.5 cm x + 5 = 125; x = 120 cm 5x = 125; x = 25 cm 25x = 125; x = 5 cm

OpenStudy (anonymous):

@Elsa213 @redeemed1312 @timeflybyegirl

OpenStudy (anonymous):

@nincompoop @Hotchellerae21 @Abhisar

OpenStudy (anonymous):

@tejasvir

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

So we know that a rectangle has two pairs of equal sides. You have to create an equation that can include that all sides will equal 125. So let's take a quick look at the options.. B. x + 5 = 125 and x = 120. This is true, yes, that 120 + 5 = 125. However, this doesn't say anything about the lengths of the rectangle. C. 5x = 125 and x = 25. Same thing here, no idea of anything about the sides of the rectangle. D. 25x = 125 and x = 5. This is basically option C. reversed. So why is the answer A.?

OpenStudy (anonymous):

It tells you 2x + 10 = 125 and that x = 57.5. The '2' in front of the x represents the two long sides of the rectangle, so that means that 57.5 is the length of one side of the rectangle.

OpenStudy (anonymous):

OMG Thanks